40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10694     Entry time: Mon Nov 10 23:14:20 2014
Author: rana 
Type: Summary 
Category: SUS 
Subject: SRM: damping gains & Optical Lever servo Tune-up 
  1.  tweaked gains for POS, PIT, YAW, SIDE by ~10-20% to get nice ringdowns with Q~5.
  2. Measured bounce rool freqs = 16.43 / 23.99 Hz. Updated the Mech Res Wiki page. Tightened up the bandstops to get back a few deg of phase. Propagated these new bandstops into the SRM SUS damping filter banks.
  3. Made and turned on a LP filters for the loops.
  4. Added a ~0.3 Hz gain boost / bubble.
  5. Set UGF to be ~2.5x below where it rings up. Estimate this to be ~3.5 Hz.
  6. First PDF shows bounce/roll peaks in OSEMs. Notice how f_roll is > sqrt(2)*f_bounce. ??
  7. Second PDF shows the OL spectra with the loops on/off. Previously there was no Boost and no LP (!!) turned on.
  8. 3rd PDF shows the modeled Bode plot of the OLG. yellow/blue is boost off/on
Attachment 1: SRM-BR.pdf  12 kB  | Hide | Hide all
SRM-BR.pdf
Attachment 2: SRM_err_141110.pdf  72 kB  | Hide | Hide all
SRM_err_141110.pdf
Attachment 3: SRM-OLG.pdf  20 kB  | Hide | Hide all
SRM-OLG.pdf
ELOG V3.1.3-